# Efficient Generation of Grids and Traversal Graphs in Compositional Spaces towards Exploration and Path Planning Exemplified in Materials

### v1

@@ 2.2 Quasi Monte Carlo
While beyond the current implementation scope of nimplex, it is beneficial to consider quasi-Monte Carlo (QMC) sampling methods, where quasi-random sequences of low discrepancy (having highly uniform coverage of all regions) are used to sample the space deterministically. Such an approach is guaranteed to be very beneficial in low-dimensional ($d\leq 3$) problems and has been implemented in thermodynamic tools, including pycalphad [44, 40] improving sampling of ternary systems. However, the QMC can become problematic as one moves to higher dimensional problems.
Firstly, the upper discrepancy bounds for QMC quickly increase with increasing $N$, unlike MC, which depends only on the number of samples; thus, MC can outperform it (thanks to better guarantees) unless a quickly (often exponentially) growing number of samples is taken (see discussion on p.271 in [45]). Because of this, even for quaternary ($d=4$) spaces, MC may be preferred for a low number of samples, even though QMC, especially with additional scrambling, can outperform it, as shown in [40].
Another significant problem in QMC is the unequal sampling of different dimensions, which can be very severe in high dimensions (see p.154 in [46]). In addition to causing under-performance in space-filling, such bias, combined with the standard alphabetical ordering of chemical components, can cause systematically worse exploration of, e.g., titanium compared to aluminum in CCMs, just based on their names.
@@ 3.1 Full
Next, one can consider the creation of a grid of uniformly distributed points, which is known to contain $\binom{d-1+n_{d}}{d-1}$ points, as discussed in Section 1.4. Similar to the random sampling discussed in Section 2, such a compositional grid cannot be constructed by simply projecting a Cartesian grid in $(N-1)$-cube as patterns will emerge (explored in detail in [40]), but it can be quickly constructed through rejecting infeasible points, as shown in Figure 5. However, it will suffer from a nearly as bad rejection rate, quantitatively dependent on both $d$ and $n_{d}$. For instance, if we consider $5\%$ spaced compositions in 9-components, the fraction of points summing to $100\%$ is $f_{M=20}(9)\approx\frac{1}{12,169}$ or $0.0082\%$.
@@
Fortunately, in their 1978 textbook, [36, ] explored the problem and gave an efficient algorithm/routine called NEXCOM to procedurally generate these simplex lattice points for arbitrary $d$ and $n_{d}$, resulting in the grid shown in Figure 5 on the right.
In the following years, several authors made various modifications to the algorithm, and the most recent one by [37, ] improves performance without sacrificing simplicity. Over the years, it has been implemented in relatively modern languages such as FORTRAN90, C, MATLAB, and Python. Now, it has been implemented in Nim language as well, with the Nim code snippet shown below.
proc simplex_grid(
dim: int, ndiv: int): Tensor[int] =
let N: int = binom(ndiv+dim-1, dim-1)
result = newTensor[int]([N, dim])
var x = zeros[int](dim)
x[dim-1] = ndiv
for j in 0..dim-1:
%**** main.tex Line 325 ****result[0, j] = x[j]
var h = dim
for i in 1..N-1:
h -= 1
let val = x[h]
x[h] = 0
x[dim-1] = val - 1
x[h-1] += 1
for j in 0..dim-1:
result[i, j] = x[j]
if val != 1:
h = dim
return result
@@ 3.2 Internal
To the best of the authors’ knowledge, something that has not been implemented before, but that is significant to exploration of CCMs (see Sec 1.2) is an algorithm to obtain only internal points of the simplex grid, i.e., points with non-zero values in all dimensions, to allow, e.g., generating all 7-component HEAs rather than all alloys in 7-component space. In principle, one can filter the output of the algorithm presented in Section 3.1; however, this may quickly become inefficient, especially for $n_{d}$ low enough as to approach $d$.
@@
We can now look at $N_{I}(d,n_{d})$ to $N_{S}(d,n_{d})$ ratio for the aforementioned case of generating all 7-component alloys. For $5\%$ grid ($n_{d}=20$) we get $\approx\frac{1}{8.5}$, and for $10\%$ grid ($n_{d}=10$) we get $\approx\frac{1}{95}$, showing a clear benefit of implementing the new method. This can be done by taking the modified-NEXCOM algorithm [37] from Section 3.1 and:
@@
proc simplex_internal_grid(
dim: int, ndiv: int): Tensor[int] =
let N: int = binom(ndiv-1, dim-1)
result = newTensor[int]([N, dim])
var x = ones[int](dim)
x[dim-1] = ndiv+1-dim
for j in 0..dim-1:
result[0, j] = x[j]
var h = dim
for i in 1..N-1:
h -= 1
let val = x[h]
%**** main.tex Line 375 ****x[h] = 1
x[dim-1] = val - 1
x[h-1] += 1
for j in 0..dim-1:
result[i, j] = x[j]
if val != 2:
h = dim
return result
@@ 4 Simplex Graph
The simplex grid algorithm presented in Section 3.1 is used commonly; however, it has an important feature that has not been utilized yet and was only briefly noted by its authors [37]. Namely, the fact that generated points are sorted in a lexicographic order (forward or reverse, depending on convention) which opens the door for using pure combinatorics for finding certain interesting relations between points at near-zero costs compared to other popular methods.
@@ 4.1 Binary
proc neighborsLink2C(i:int, x:Tensor,
neighbors: var seq[seq[int]]): void =
if x[0] != 0:
neighbors[i].add(i+1)
if x[1] != 0:
neighbors[i].add(i-1)
@@ 4.2 Ternary
proc neighborsLink3C(…,
ndiv: int): void =
let jump0 = 1
let jump1 = 1+ndiv-x[0]
if x[0] != 0:
neighbors[i].add(i-jump1)
neighbors[i].add(i-jump1-jump0)
if x[1] != 0:
neighbors[i].add(i-jump0)
neighbors[i].add(i+jump1-jump0)
if x[2] != 0:
neighbors[i].add(i+jump0)
neighbors[i].add(i+jump1)
@@ 4.3 N-Dimensional
proc neighborsLink(…): void =
var jumps = newSeq[int](dim-1)
%**** main.tex Line 475 ****jumps[0] = 1 #binom(a,0)=1
for j in 1..<(dim-1):
jumps[j] = binom(
j+ndiv-sum(x[0..(dim-2-j)]), j)
var trans: int
for order in 0..(dim-2):
trans = 0
if x[order] != 0:
for dir in 0..(dim-2-order):
temp += jumps[dim-2-order-dir]
neighbors[i].add(i - trans)
neighbors[i - trans].add(i)
@@ 4.4 Simplex Graph Complexes
Suppose one tries to connect elemental compositions A and F, but assumes prior knowledge that they cannot be combined directly in any quantity, and also knows that (1) A is compatible with B and C, (2) F is compatible with D and E, but (3) B and E are incompatible in any quantity, (4) C and D are incompatible in any quantity. Furthermore, (5) G and H are individually compatible with B and D, and (6) I and J are individually compatible with C and E.
These rules can be used to set up a problem graph like in the top of Figure 9, encoding everything that is known about the system a priori and limiting the solution space from full $\binom{10-1+12}{12}\approx 300,000$ to 2×($3122\times\binom{3-1+12}{12}+10\times\binom{2-1+12}{12}=$ 312, or three orders of magnitude.
@@
One can quickly demonstrate the benefits of such ability by looking at the SS316 to Ti-6Al-4V problem studied by [22, ]. After idealizing and anonymizing the components, it becomes a problem where one tries to combine compositions A with G, which cannot be combined directly in almost any quantity, and also knows that (1) system ABC is highly feasible across it, but (2) C cannot be combined directly with G in any quantity, and (3) a complete path from pure B to G is not possible. In this case, a simple problem setup is to look at several BC? and BG? pairs, forming parallel pathways from ABC to G. This is depicted in Figure 10 for 3 candidates D, E, F, forming 6 ternary spaces to consider, but nothing limits the method to be extended to an arbitrary number of candidates while still retaining its linear complexity.
@@ 4.5 Discussion of Exploration
This, in turn, enables one to harvest many general-purpose graph traversal algorithms, which are actively researched and available through high-performance libraries. For instance, to navigate against constraints, the $A^{*}$ algorithm [47] can be used with such a heuristic and is mathematically guaranteed to find the shortest feasible compositional path while exploring the least number of nodes [48], what can be critical if the shortest path is necessary while each evaluation is relatively expensive. Then, if one tries to find a feasible solution first and then improve on it, modifications of the $A^{*}$ algorithms such as the $RWA^{*}$ [49] can be used to first make it more greedy and then gradually move towards $A^{*}$ to obtain the optimal solution if sufficient resources are available. Alternatively, for highly complex problems where exploration needs to proceed quickly towards the goal but the optimality guarantees are not needed, one can use a search algorithm based on the Monte Carlo tree search (MCTS), which has been famously used in conjunction with an ML model to master the game of Go [50].
5 Summary and Conclusion
@@
    In Section 3.2, a new algorithm was developed to efficiently create internal (subspace-exclusive) grids in simplex spaces based on an algorithm from the literature (modified-NEXCOM [37]). It is beneficial to performance in cases of, for instance, sampling only $d$-component materials in $d$-component chemical space without considering lower-order points.
@@
The nimplex software described in this work has been published as free open-source software (FOSS) under the MIT license. It can be effortlessly used as a native Nim library, native Python library, or Command Line Interface (CLI) tool interfacing with nearly any language through binary data or plain text.
@@ Acknowledgments
We would also like to thank Luke Myers and Ricardo Amaral for testing code exercises and proofreading the documentation.
@@ Appendix A Appendix A
proc neighborsLink4C(…): void =
let jump0 = 1 #binom(x, 0)=1
let jump1 = binom(1+ndiv-x[0]-x[1], 1)
let jump2 = binom(2+ndiv-x[0], 2)
[…]
# quaternary
neighbors[i].add(i - jump2)
# quaternary
neighbors[i].add(i - jump2 - jump1)
# quaternary
neighbors[i].add(i - jump2 - jump1 - jump0)
[…]
# ternary
neighbors[i].add(i - jump1)
# ternary
neighbors[i].add(i - jump1 - jump0)
# quaternary
neighbors[i].add(i + jump2 - jump1 - x[1])
[…]
# binary
%**** main.tex Line 675 ****neighbors[i].add(i - jump0)
# ternary
neighbors[i].add(i + jump1 - jump0)
# quaternary
neighbors[i].add(i + jump2 - jump0 - x[1])
[…]
# binary
neighbors[i].add(i + jump0)
# ternary
neighbors[i].add(i + jump1)
# quaternary
neighbors[i].add(i + jump2 - x[1])

### v2

@@ 2.2 Quasi Monte Carlo
While beyond the current implementation scope of nimplex, it is beneficial to consider quasi-Monte Carlo (QMC) sampling methods, where quasi-random sequences of low discrepancy (having highly uniform coverage of all regions) are used to sample the space deterministically. Such an approach is guaranteed to be very beneficial in low-dimensional ($d\leq 3$) problems and has been implemented in thermodynamic tools, including pycalphad [Otis2017Pycalphad:Python, Otis2017AnSystems] improving sampling of ternary systems. However, the QMC can become problematic as one moves to higher dimensional problems.
Firstly, the upper discrepancy bounds for QMC quickly increase with increasing $N$, unlike MC, which depends only on the number of samples; thus, MC can outperform it (thanks to better guarantees) unless a quickly (often exponentially) growing number of samples is taken (see discussion on p.271 in [Asmussen2007StochasticAnalysis]). Because of this, even for quaternary ($d=4$) spaces, MC may be preferred for a low number of samples, even though QMC, especially with additional scrambling, can outperform it, as shown in [Otis2017AnSystems].
Another significant problem in QMC is the unequal sampling of different dimensions, which can be very severe in high dimensions (see p.154 in [Lemieux2009MonteSampling]). In addition to causing under-performance in space-filling, such bias, combined with the standard alphabetical ordering of chemical components, can cause systematically worse exploration of, e.g., titanium compared to aluminum in CCMs, just based on their names.
@@ 3.1 Full
[…]
@@
[…]
{minted}
[linenos=true, mathescape]nim proc simplex_grid( dim: int, ndiv: int): Tensor[int] = let N: int = binom(ndiv+dim-1, dim-1) result = newTensor[int]([N, dim]) var x = zeros[int](dim) x[dim-1] = ndiv for j in 0..dim-1: result[0, j] = x[j] var h = dim for i in 1..N-1: h -= 1 let val = x[h] x[h] = 0 x[dim-1] = val - 1 x[h-1] += 1 for j in 0..dim-1: result[i, j] = x[j] if val != 1: h = dim return result
@@ 3.2 Internal
To the best of the authors’ knowledge, a minor modification that has not been explicitly implemented before, but that is significant to exploration of CCMs (see Sec 1.2) is an algorithm to obtain only internal points of the simplex grid, i.e., points with non-zero values in all dimensions, to allow, e.g., generating all 7-component HEAs rather than all alloys in 7-component space. In principle, one can filter the output of the algorithm presented in Section 3.1; however, this may quickly become inefficient, especially for $n_{d}$ low enough as to approach $d$.
@@
We can now look at $N_{I}(d,n_{d})$ to $N_{S}(d,n_{d})$ ratio for the aforementioned case of generating all 7-component alloys. For $5\%$ grid ($n_{d}=20$) we get $\approx\frac{1}{8.5}$, and for $10\%$ grid ($n_{d}=10$) we get $\approx\frac{1}{95}$, showing a clear benefit of implementing the new method. This is often done by using the full grid algorithm but with $d$ subtracted from $n_{d}$, followed by the addition of $1$ to all grid elements, but it can also be accomplished by a dedicated algorithm to avoid any performance overhead. In this work, the latter is done by taking the modified-NEXCOM algorithm [Chasalow1995AlgorithmPoints] from Section 3.1 and:
@@
{minted}
[linenos=true, mathescape]nim proc simplex_internal_grid( dim: int, ndiv: int): Tensor[int] = let N: int = binom(ndiv-1, dim-1) result = newTensor[int]([N, dim]) var x = ones[int](dim) x[dim-1] = ndiv+1-dim for j in 0..dim-1: result[0, j] = x[j] var h = dim for i in 1..N-1: h -= 1 let val = x[h] x[h] = 1 x[dim-1] = val - 1 x[h-1] += 1 for j in 0..dim-1: result[i, j] = x[j] if val != 2: h = dim return result
@@ 4 Simplex Graph
The simplex grid algorithm presented in Section 3.1 is used commonly; however, it has an important feature that has not been utilized yet and was only briefly noted by its authors [Chasalow1995AlgorithmPoints]. Namely, the fact that generated points are sorted in a lexicographic order (forward or reverse, depending on convention) which opens the door for using pure combinatorics for finding certain interesting relations between points at near-zero costs compared to other popular methods.
@@ 4.1 Binary
{minted}
[linenos=true, mathescape=true]nim proc neighborsLink2C(i:int, x:Tensor, neighbors: var seq[seq[int]]): void = if x[0] != 0: neighbors[i].add(i+1) if x[1] != 0: neighbors[i].add(i-1)
@@ 4.2 Ternary
{minted}
[linenos=true, mathescape=true]nim proc neighborsLink3C(…, ndiv: int): void = let jump0 = 1 let jump1 = 1+ndiv-x[0] if x[0] != 0: neighbors[i].add(i-jump1) neighbors[i].add(i-jump1-jump0) if x[1] != 0: neighbors[i].add(i-jump0) neighbors[i].add(i+jump1-jump0) if x[2] != 0: neighbors[i].add(i+jump0) neighbors[i].add(i+jump1)
@@ 4.3 N-Dimensional
{minted}
[linenos=true, mathescape=true]nim proc neighborsLink(…): void = var jumps = newSeq[int](dim-1) jumps[0] = 1 #binom(a,0)=1 for j in 1..<(dim-1): jumps[j] = binom( j+ndiv-sum(x[0..(dim-2-j)]), j) var trans: int for order in 0..(dim-2): trans = 0 if x[order] != 0: for dir in 0..(dim-2-order): temp += jumps[dim-2-order-dir] neighbors[i].add(i - trans) neighbors[i - trans].add(i)
@@ 4.4 Simplex Graph Complexes
Suppose one tries to connect elemental compositions A and F, but assumes prior knowledge that they cannot be combined directly in any quantity, and also knows that (1) A is compatible with B and C, (2) F is compatible with D and E, but (3) B and E are incompatible in any quantity, (4) C and D are incompatible in any quantity. Furthermore, (5) G and H are individually compatible with B and D, and (6) I and J are individually compatible with C and E. These rules can be used to set up a problem graph like in the top of Figure 9, encoding everything that is known about the system a priori and limiting the solution space from full $\binom{10-1+12}{12}\approx 300,000$ to 2×($3122\times\binom{3-1+12}{12}+10\times\binom{2-1+12}{12}=$ 312, or three orders of magnitude.
@@ 4.5 Exploration
This, in turn, enables one to harvest many general-purpose graph traversal algorithms, which are actively researched and available through high-performance libraries. For instance, to navigate against constraints, the $A^{*}$ algorithm [Hart1968APaths] can be used with such a heuristic and is mathematically guaranteed to find the shortest feasible compositional path while exploring the least number of nodes [Dechter1985GeneralizedA], what can be critical if the shortest path is necessary while each evaluation is relatively expensive. Then, if one tries to find a feasible solution first and then improve on it, modifications of the $A^{*}$ algorithms such as the $RWA^{*}$ [Bhatia2021OnA] can be used to first make it more greedy and then gradually move towards $A^{*}$ to obtain the optimal solution if sufficient resources are available. Alternatively, for highly complex problems where exploration needs to proceed quickly towards the goal but the optimality guarantees are not needed, one can use a search algorithm based on the Monte Carlo tree search (MCTS), which has been famously used in conjunction with an ML model to master the game of Go [Silver2016MasteringSearch].
5 Discussion
@@
    In Section 3.2, a new algorithm was developed to efficiently create internal (subspace-exclusive) grids in simplex spaces based on an algorithm from the literature (modified-NEXCOM [Chasalow1995AlgorithmPoints]) without introducing any overhead. It is beneficial in cases of, for instance, sampling only $d$-component materials in $d$-component chemical space without considering lower-order points.
@@
Data Availability
No datasets were generated or analyzed during the current study. The research presents a theoretical algorithm and mathematical equations that do not rely on empirical data.
@@
The nimplex software described in this work has been published as free open-source software (FOSS) under the MIT license. It can be effortlessly used as a native Nim library, native compiled Python library, or Command Line Interface (CLI) tool interfacing with nearly any language through binary data or plain text.
@@ Acknowledgments
We would also like to thank Luke Myers, Ricardo Amaral, and Alexander Richter for testing code exercises and proofreading the documentation.
Author Contributions
Competing Interests
The authors declare no competing interests.
[…]
@@ Appendix A Appendix A
{minted}
nim proc neighborsLink4C(…): void = let jump0 = 1 #binom(x, 0)=1 let jump1 = binom(1+ndiv-x[0]-x[1], 1) let jump2 = binom(2+ndiv-x[0], 2)
if x[0] != 0: # quaternary neighbors[i].add(i - jump2) # quaternary neighbors[i].add(i - jump2 - jump1) # quaternary neighbors[i].add(i - jump2 - jump1 - jump0)
if x[1] != 0: # ternary neighbors[i].add(i - jump1) # ternary neighbors[i].add(i - jump1 - jump0) # quaternary neighbors[i].add(i + jump2 - jump1 - x[1])
if x[2] != 0: # binary neighbors[i].add(i - jump0) # ternary neighbors[i].add(i + jump1 - jump0) # quaternary neighbors[i].add(i + jump2 - jump0 - x[1])
if x[3] != 0: # binary neighbors[i].add(i + jump0) # ternary neighbors[i].add(i + jump1) # quaternary neighbors[i].add(i + jump2 - x[1])